Please give explanations or show the steps. Thank you.

6. What is angle of an intradermal injection?

7. Order: Atropine 0.8 mg IV pre-medication 30 minutes before anesthesia, in stock 400mcg/mL.

How many mL needed for IV?

8. Round up to nearest hundredth: 7.138

9. Convert 70 inches to centimeter.

10. Provider orders KCL elixir 60mEq PO daily, in stock you have KCL elixir 40mEq per 15mL.

How many mL should client receive?
Answer #1

Ans:- 6.

5°-15°,

Angle for an intradermal injection is should be between 5 - 15 degree.

7.

Ordered atropine is 0.8mg ,

Available atropine is 400mcg/ ml or 0.4mg / ml.

So, there we the need of 2 ml for IV to administer 0.8mg.

8.

7.14.

Round up to nearest hundredth of 7.138 will be 7.14 because in the place of thousands their 8 that is more than 5 then in this case the digit which is at the place of hundred will become 4 with the addition of 1, if the digit at thousands would be less then 5 then the digit at the place of hundred will remain same.

9.

177.8 cm,

70 inches to centimeter will be 177.8 cm because

1 inch is equal to 2.54 cm ,

Then 70 cm = 2.5×70= 177.8 cm

10.

22.5 ml,

Ordered KCL elixir is 60 mEq,

In stock available KCL elixir 40 mEq/ 15 ml,

Then ,

1 mEq = 15 / 40,

In 60 mEq = 15/40×60 = 22.5 ml